Solve for Y(s), the Laplace transform of the solution y(t) to the initial value problem below. y +By = 42 -9. y(O)=0. y(0)
Answer #1

The Given equation is - y + 6y = 479; 4(o)=0 and y(0) = -2 Taking <saplaee transform both sides weget L {y } + L {64} = L {$184+6) S3 2 8-95² A. B. 53152+6) A(8+6) + B.s? S3 (5²76) (A+B) s? + GA. 53 15+6) Comparing; A+B = -9 and 6A = 8 8/6 on. S +soi S 2 1 + Y(s) 31 1 18. s To 31 18 S3 s²6 5) (s’+6 Taking Inverse Looplace transform weget yGt) = } L Lola 2 s²+6 in t2 = 3